OpenStudy (anonymous):

x 7x 2 - + - < - The solution set is? 5 10 5

OpenStudy (tyteen4a03):

\(\frac{x}{5} + \frac{7x}{10} < \frac{2}{5}\) Firstly find the GCD of the three fractions, then multiply all variables by the GCD to get an inequality with no fractions. Then continue solving.